OpenStudy (anonymous):

solve using the elimination method. If the system has no solution or an infinite number of the solutions state this. -2x + 9y = 56, 9x -9y =0

OpenStudy (anonymous):

-2x + 9y = 56, 9x -9y=0 In elimination, you basically put both equations into one big one, with everything on the left side of equal sign together and everything on the right together. -2x + 9y +9x - 9y = 56 + 0 Now, collect like terms 7x = 56 (+9y-9y cancel each other out) Solve for x. 56/7 = 8. x=8. Now to solve for Y, just put x into one of your given equations 9x - 9y = 0 9(8) - 9y=0 (move 9y over ) 72 = 9y 8=y